Define drift velocity and derive an expression for it

Grade:12

1 Answers

Arun
6 years ago
Dreift velocity is the velocity attained by any particle such as electron when kept in an electric field
Let the electron have charge e,  mass m and kept in a electric field E inside a conductor.
The force experienced by the electron F = eE
Acceleration due to the force   = eE/m
Now displacement s due to the acceleration s = ½at2
Average velocity  vd  (drift velocity = Total displacement / total time taken
= s/t
vd = eE/2m